# Are We Overlooking the Dimensions? Learning Latent Hierarchical Channel Structure for High-Dimensional Time Series Forecasting

### v1

@@ I.5 Evaluation Metrics
With a typical choice $r=16$, UCast lowers both time and memory by a factor of $16$ compared with a flat channel Transformer, yet still retains the expressive power of attention through its latent-query hierarchy.
Appendix K Dataset Description
Temp (Global Temp) and Wind (Global Wind). We adapt this dataset from Corrformer [50]. Global Temp&Wind dataset is from the National Centers for Environmental Information (NCEI)  [^2] https://www.ncei.noaa.gov/. This dataset contains the hourly averaged wind speed and hourly temperature of 3,850 stations (7,700 dimensions) around the world from 01/01/2019 to 12/31/2020, 17,544 timesteps in total. This dataset is well established. However, we find there are some constant columns in the datasets, which means the temp or wind speed never change. We remove these columns.
Solar (NREL Solar Power). NREL Solar Power is collected from National Renewable Energy Laboratory (NREL)  [^3] https://www.nrel.gov/grid/solar-power-data.html. One of the most popular datasets for time series forecasting is Solar-Energy [20], which is sourced from NREL. However, it includes data from only 137 solar photovoltaic (PV) plants in Alabama, limiting the dataset to 137 dimensions. To increase the dimensionality, we collect data from 5,166 PV plants across 47 U.S. states, excluding Alaska, Hawaii, and North Dakota, using data from NREL. The data source is clean so we don’t conduct further preprocessing on it.
Traffic-CA (LargeST-CA), Traffic-GLA (LargeST-GLA), Traffic-GBA (LargeST-GBA). The source of this dataset is the Caltrans Performance Measurement System (PeMS) [^4] https://pems.dot.ca.gov/, and we utilize its preprocessed version[24]. LargeST-CA encompassing a total of 8,600 sensors in California and each sensor contains five years of traffic flow data (from 2017 to 2021) with a 5-minute interval, resulting in a total of 525,888 time frames. LargeST-CA consists of three subsets: (1) LargeST-GLA includes 3,834 sensors installed across five counties in the Greater Los Angeles (GLA) area: Los Angeles, Orange, Riverside, San Bernardino, and Ventura. (2) LargeST-GBA contains 2,352 sensors distributed across 11 counties in the Greater Bay Area (GBA): Alameda, Contra Costa, Marin, Napa, San Benito, San Francisco, San Mateo, Santa Clara, Santa Cruz, Solano, and Sonoma. (3) LargeST-SD, the smallest subset, consists of 716 sensors located exclusively in San Diego County. Since LargeST-SD has only 716 dimensions, we do not include it in our benchmark. For the 525,888 original time frames, we aggregate the data into 1-hour intervals, resulting in 43,824 time frames. We then remove sensors with more than 100 missing (NaN) values, and apply linear interpolation to fill the remaining missing values.
Neurolib. Simulations made use of the Neurolib code [4]. We utilize the brain structure in ABIDE dataset [9].
SIRS. Simulations extend the classical SIR model [3] by incorporating temporary immunity, resulting in dynamic transitions among Susceptible, Infected, and Recovered groups within a single region. To enhance realism, parameter noise, and seasonal variation are included. For multiple regions, each compartment has a fixed transfer rate between regions to simulate commuting and spatial transmission dynamics.
Air Quality (Chinese Air Quality). This dataset includes 7 air quality indicators: $AQI$, $CO$, $NO_{2}$, $O_{3}$, $PM10$, $PM2.5$, and $SO_{2}$ of 786 distinct cities in China. We remove dimensions with more than 100 missing values and interpolate other dimensions linearly, resulting in 1,106 dimensions.
M5. We adapt this dataset from a Kaggle competition [18]. It covers stores in three U.S. states, i.e. California, Texas, and Wisconsin, and includes item-level sales data, department and product category details, and store information. The dataset records daily sales for each item from 07/01/2015 to 06/30/2022 (a total of 1,947 days) across different departments and stores, resulting in 30,490 time series. Due to its sparsity (i.e., a large number of zero values), we aggregate sales by summing each item’s sales across all departments and stores, reducing the dataset to 3,049 aggregated items.
Measles (Measles England). The measles dataset [31, 28, 27], contains biweekly measles infections from 387 regions across England and Wales from 1944–1965. For each region, the dataset includes three features: inferred infections, population, and suspected cases, resulting in 1,161 dimensions.
Wiki-20k (Wikipedia Web Traffic) The dataset is adapted from the Monash Time Series Forecasting Repository  [^8] https://zenodo.org/records/7370977. It comprises 145,063 time series representing web traffic for various Wikipedia pages from July 1, 2015, to June 30, 2022 (2,557 days). This dataset is an extended version of the one used in the Kaggle Wikipedia Web Traffic Forecasting Competition  [^9] https://www.kaggle.com/c/web-traffic-time-series-forecasting/data. Due to data sparsity, we remove the Wikipedia pages containing NaN values, resulting in a final dataset of 112,333 Wikipedia pages, each corresponding to one dimension. Due to the original dataset’s dimension is too large and is not correlated, we select the first 20k dimensions to construct Wiki-20k dataset for evaluation.
Mobility (Google Community Mobility). The Community Mobility dataset [15] aims to assist in analyzing the impact of COVID-19 on community movement patterns. The dataset covers movement trends over time (from February 15, 2020, to October 15, 2022, spanning 974 days) across 4,334 regions in different countries. It categorizes mobility data into six place types: retail and recreation (restaurants, cafes, shopping centers, theme parks, museums, libraries, and movie theaters), groceries and pharmacies (grocery stores, food warehouses, farmers markets, specialty food shops, drug stores, and pharmacies), parks (local and national parks, public beaches, marinas, dog parks, plazas, and public gardens), transit stations (subway, bus, and train stations), workplaces (places of employment), and residential (places of residence). The dataset initially consists of 26,004 dimensions, derived from the six place categories across 4,334 regions. However, due to data sparsity, we exclude regions containing NaN values, reducing the dataset to 971 regions and 5,826 dimensions. The values in the dataset represent changes in visit frequency and duration of stay at various locations relative to the baseline (i.e. the median value for the same day of the week during the five-week period from January 3 to February 6, 2020.). These changes are computed using aggregated and anonymized data, similar to the methodology employed in Google Maps to estimate popular visit times.
Appendix L Channel Correlation Statistic
We follow TFB [38] to compute channel correlation. In multivariate time series, different variables often share common temporal patterns due to underlying shared factors. To quantify this, we extract 22 features per channel using Catch22 [29]. These features serve as fixed-length representations of each time series channel. We then compute the pairwise Pearson correlation coefficients [8] between all channel representations and report the mean and variance. The results are shown in the “Correlation” column of Table 2. All datasets exhibit high average channel correlation, confirming the presence of strong inter-variable dependencies in these high-dimensional time series.
Appendix M Implementation Details
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following [25, 30, 44, 48, 49, 58, 53]:
Experiment Details. All experiments are implemented in PyTorch [37] and conducted on a cluster equipped with 8 NVIDIA A100 GPUs. The baselines are trained using their default configurations as reported in their respective papers, and with further hyperparameter searching (Learning Rate in {0.01, 0.001, 0.0001}, Input Sequence Length $T$ in $3\times S$, $4\times S$, $5\times S$}), where $S$ is prediction length. ADAM optimizer [19] is used with MSE loss for the training of all the models. We apply early stopping with a patience value of 5 epochs. The batch size is initially set to 32. If an out-of-memory (OOM) error occurs, the batch size is automatically halved until the issue is resolved. For U-Cast, we default to using non-stationary [26] normalization for all datasets. We default to set the batch size to 128, the number of layers $L$ to 2, the hidden dimension $d$ to 512, and the reduction ratio $r$ to 16 for all the datasets. Additional detailed U-Cast model configuration information is presented in Table 6. We conduct hyperparameter sensitivity analysis in Appendix N.
Table 6: U-Cast hyperparameters used for each dataset.
Dataset Learning Rate Input Sequence Length $T$ $\alpha$
Atec 0.002 $3\times S$ 0.001
Air Quality 0.0005 $4\times S$ 1
Temp 0.0005 $3\times S$ 0.01
Wind 0.001 $4\times S$ 0.001
Mobility 0.002 $4\times S$ 10
Traffic-CA 0.001 $4\times S$ 1
Traffic-GBA 0.001 $4\times S$ 0.01
Traffic-GLA 0.001 $4\times S$ 1
M5 0.0005 $4\times S$ 0.1
Measles 0.0005 $3\times S$ 0.001
Neurolib 0.002 $4\times S$ 0.001
Solar 0.0005 $4\times S$ 0.001
SIRS 0.0005 $4\times S$ 1
Meters 0.0005 $4\times S$ 0.001
SP500 0.001 $3\times S$ 1
Wiki-20k 0.0005 $4\times S$ 10
Appendix N Hyperparameter Sensitivity Analysis
@@
Figure 5: Performance (MSE) on China Air Quality dataset under different hyperparameters.
@@
Sensitivity Analysis of the full-rank regularization strength $\alpha$. We analyze how the regularization coefficient $\alpha$ affects the performance of U-Cast by varying it across several orders of magnitude. As shown in Figure 5 (a), the model achieves the best performance when $\alpha$ is set to a small value (0.001–0.01). As $\alpha$ increases, the forecasting error (MSE) gradually rises, suggesting that overly strong regularization may overconstrain the channel representations, thereby reducing the model’s ability to capture informative structures. Based on this analysis, we select $\alpha=0.01$ for a good balance between disentanglement and flexibility.
@@
Sensitivity Analysis of the number of hierarchical layers $L$. We evaluate the effect of the number of hierarchical latent query layers $L$ on forecasting performance, where a lower mean squared error (MSE) indicates better accuracy. As shown in Figure 5 (b), increasing $L$ from 2 to 3 reduces the MSE, indicating worse performance. This suggests that deeper hierarchies may not capture more useful structure and could lead to redundancy or overfitting. Based on this analysis, we set $L=2$ in our main experiments.
@@
Sensitivity Analysis of the reduction ratio $r$. We investigate how the reduction ratio $r$, which determines the compression of channel representations at each layer, influences performance. As shown in Figure 5 (c), the model achieves relatively stable performance across a wide range of $r$ values. However, extremely large or small values can lead to slight degradation. For instance, very low reduction (e.g., $r=2$) leads to higher computational cost without performance gain, while aggressive reduction (e.g., $r=32$) may discard useful information. We find that $r=16$ offers a favorable trade-off between efficiency and accuracy.
Figure 6: Training time across datasets.
Appendix O Efficiency
We compare the training time per batch with recent state-of-the-art models in Figure 6. The x-axis represents different datasets, arranged from left to right in order of increasing dimensionality. The results show that: (1) U-Cast consistently achieves favorable training efficiency across varying dimensionalities; (2) As dimensionality increases, the efficiency advantage of U-Cast over other non-linear methods becomes more pronounced; (3) The efficiency of CI-based methods such as PAttn and PatchTST is affected by both the channel and temporal dimensions. When the channel dimensionality becomes very large, as in the Wiki20000 dataset, its efficiency degrades significantly; (4) The efficiency of CD-based methods, including iTransformer, CCM, and U-Cast, is primarily influenced by the number of input channel dimensions.

### v2

@@ Appendix I Theorem 3 Proof
With a typical choice $r=16$, UCast lowers both time and memory by a factor of $16$ compared with a flat channel Transformer, yet still retains the expressive power of attention through its latent‑query hierarchy.
Appendix M Empirical Evidence of Hierarchical Structures
Table 9: CPCC values of hierarchical clustering applied to real-world datasets. Higher values (closer to 11) indicate stronger hierarchical structure.
Dataset M5 Meter Temp Wind Atec Mobility SP500 Air Quality Measles
CPCC 0.7852 0.7705 0.9057 0.7478 0.8866 0.8338 0.9484 0.7429 0.8950
To empirically evaluate the hierarchical structure of real-world datasets, we apply hierarchical clustering to the data and assess the resulting structure using the cophenetic correlation coefficient (CPCC) (Saraçli et al., 2013). A CPCC value closer to $1$ indicates that the dendrogram generated by hierarchical clustering preserves the original pairwise distances between data points more accurately (Kumar & Toshniwal, 2016), which implies that the features in the dataset exhibit a stronger hierarchical structure.
Appendix N Empirical Validation for Channel Dependency Findings on Real-World Data
Table 10: MAE performance of CI model DLinear and CD model U-Cast on Wiki datasets with different channel subsets.
Channels DLinear U-Cast
Wiki-20k 20,000 0.394 0.302
Wiki-10k 10,000 0.489 0.385
Wiki-2k 2,000 0.669 0.659
Wiki-0.2k 200 0.697 0.782
To bridge the synthetic analysis to a real-world setting, we conducted an additional experiment. To reduce the effect of dataset-specific bias, we fixed the dataset to Wiki-20k, which is a real-world dataset, and randomly selected different numbers of channels to construct Wiki-10k, Wiki-2k, and Wiki-0.2k. We then evaluated the MAE performance of the CI model DLinear and the CD model U-Cast on these datasets. As shown in Table 10, CI models perform better when the dimensionality is low, but as the dimensionality increases, CD models gradually outperform CI models. This observation is consistent with the conclusion in Section 3.2.
Appendix O Hyperparameter Sensitivity Analysis
@@
Figure 5: Performance (MSE) on Air Quality dataset under different hyperparameters.
Figure 6: Comparison of loss scales between the MSE loss $\mathcal{L}_{mse}$ and the full-rank regularization loss $\mathcal{L}_{cov}$. Left: Atec, Right: Temp.
@@
Sensitivity Analysis of Full-Rank Regularization Strength $\alpha$. We investigate how the regularization coefficient $\alpha$ influences the performance of U-Cast by varying it across several orders of magnitude. As shown in Figure 6 (a), the best performance is achieved when $\alpha$ is set to a small value (0.001–0.01). This behavior can be explained by the imbalance in the scales of the two losses: the full-rank regularization loss $\mathcal{L}_{cov}$ is roughly an order of magnitude larger than the MSE loss $\mathcal{L}_{mse}$ (see Figure 6). If $\alpha$ is too large, the optimization is dominated by $\mathcal{L}_{cov}$, which may overconstrain the latent channel representations and reduce the model’s ability to capture informative structures. Using a smaller $\alpha$ helps balance the two objectives, stabilizes gradient updates, and leads to better forecasting accuracy. Based on this analysis, we set $\alpha=0.01$ to achieve a good trade-off between disentanglement and predictive performance.
@@
Sensitivity Analysis of Number of Hierarchical Layers $L$. We evaluate the effect of the number of hierarchical latent query layers $L$ on forecasting performance, where a lower mean squared error (MSE) indicates better accuracy. As shown in Figure 5 (b), increasing $L$ from 2 to 3 reduces the MSE, indicating worse performance. This suggests that deeper hierarchies may not capture more useful structure and could lead to redundancy or overfitting. Based on this analysis, we use a unified setting of layer number $L=2$ across all 16 datasets from different domains, which consistently achieves a stable trade-off between accuracy and efficiency without the need for per-dataset hyperparameter tuning.
@@
Sensitivity Analysis of Reduction Ratio $r$. We investigate how the reduction ratio $r$, which determines the compression of channel representations at each layer, influences performance. As shown in Figure 5 (c), the model achieves relatively stable performance across a wide range of $r$ values. However, extremely large or small values can lead to slight degradation. For instance, very low reduction (e.g., $r=2$) leads to higher computational cost without performance gain, while aggressive reduction (e.g., $r=32$) may discard useful information. We find that $r=16$ offers a favorable trade-off between efficiency and accuracy. Thus, we use a unified setting of layer number $r=16$ across all 16 datasets.
Figure 7: Comparison of performance (MSE) among U-Cast, PatchTST, and iTransformer on three datasets (Air Quality, M5, SIRS) under varying input lengths $T$.
Sensitivity Analysis of Input Length $T$. To directly assess the impact of input length and to further validate the findings from our Time-HD benchmark, we conduct an extended experiment. We evaluate PatchTST, iTransformer, and U-Cast on three datasets across input lengths set to $2\times$, $3\times$, and $4\times$ the prediction horizon ($S$). The results are presented in Figure 7. These experiments lead to three main observations. First, longer input sequences generally improve performance in both CI and CD models on most datasets. Second, the effect varies by dataset. For instance, in the SIRS dataset, all models achieve their best results with the shortest input length, likely because its short-term epidemiological dynamics make longer histories add noise. This variability highlights the usefulness of Time-HD as a benchmark for testing robustness under diverse temporal structures. Third, U-Cast remains competitive or superior across most datasets and horizons, and CD models (U-Cast and iTransformer) as a group continue to perform strongly when cross-channel structure provides informative signals.
@@
Appendix Q Efficiency
We compare the training time per batch with recent state-of-the-art models in Figure 8 and Figure 9. In Figure 9, the Atec, Temp, and Wind datasets are removed to provide a clearer view of the efficiency trend at higher dimensions. The x-axis corresponds to different datasets, arranged from left to right in increasing order of dimensionality.
The results indicate that: (1) U-Cast consistently maintains favorable training efficiency across different dimensionalities; (2) Its efficiency advantage over other non-linear methods grows as dimensionality increases; (3) The efficiency of CI-based methods (e.g., PAttn, PatchTST) is sensitive to both channel and temporal dimensions. In Figure 9, when channel dimensionality is very large, as in Wiki-20k, their efficiency drops sharply; similarly, in Figure 8, longer input lengths (Atec, Temp, Wind) also lead to significant degradation; (4) As shown in Figure 9, the efficiency of CD-based methods (iTransformer, U-Cast) is mainly affected by the number of input channels. However, U-Cast is more robust than iTransformer, showing more stable efficiency as dimensionality increases.
The detailed efficiency comparison of both the training and inference phases is provided in Appendix T. In both stages, U-Cast achieves the best trade-off between performance and efficiency.
Figure 8: Training time across datasets. The x-axis represents different datasets, arranged from left to right in order of increasing dimensionality.
Figure 9: Training time across datasets (w/o Atec, Temp, and Wind datasets). The x-axis represents different datasets, arranged from left to right in order of increasing dimensionality.
Appendix R Additional Baselines
Table 12: Preliminary forecasting results of MCANN and NECPlus on selected datasets. Lower values indicate better performance.
MCANN (MSE) MCANN (MAE) NECPlus (MSE) NECPlus (MAE)
Measles 1.727 0.851 1.160 0.722
Air Quality 0.466 0.457 0.447 0.449
Mobility 3.471 1.367 1.815 0.899
M5 17.031 1.441 16.842 1.387
We additionally include two methods, MCANN Li & Anastasiu (2025) and NECPlus Li et al. (2023), which further broaden the coverage of our benchmark. Both methods have been integrated into our Time-HD-Lib. For NECPlus, we adopt the standard model implementation. Preliminary results are presented in Table 12.
Appendix S Showcases
For clear comparison, we present test set showcases in Appendix U, where U-Cast shows better performance.
Appendix T Detailed Efficiency Comparison
Appendix U Showcases
Figure 10: Prediction cases from Atec by different models under the input-S-predict-S settings.
Figure 11: Prediction cases from Temp by different models under the input-S-predict-S settings.
Figure 12: Prediction cases from Wind by different models under the input-S-predict-S settings.
Figure 13: Prediction cases from Mobility by different models under the input-S-predict-S settings.
Figure 14: Prediction cases from Measles by different models under the input-S-predict-S settings.
Figure 15: Prediction cases from SIRS by different models under the input-S-predict-S settings.